

********************************************************************
UARS WEEKLY STATUS REPORT - 20 July 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation scan
data this past week without apparent problems.
The SUSIM team was saddened by the death of Dr. Guenter Brueckner,
Principal Investigator, on July 11, 1998. His enthusiasm for solar physics
was enormous and remained to the end. We will miss him as a colleague
and as a friend.
***********
PEM
***********
PEM was not active.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS has been operated in a very low power short duration diagnostic mode to
obtain engineering data on the scan mechanism. Additional tests are being
conducted to determine the optimal operating mode for the upcoming south
viewing observation period.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 04 July 1998 (Orbit 37217, GMT Day 185) through 10 July
1998 (Orbit 37321, GMT Day 191).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 189/1814 37288 Scan Test
HRDI 190/2255 37306 Off
SUSIM 192/2304 37336 Off
MLS 193/1008 37343 Scan Test
A summary of Instrument operations supplied by the Mission Planning Group
for the month of June 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle is now decreasing toward a minimum angle of 0.0 degrees on
15 July 1998 (DOY 196).
Two "manually generated" SSPP stellar offset experiments were performed on
GMT Days 188 and 190 as requested by the SOLSTICE operations team.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.5 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 03 July 1998 and 10 July 1998 are provided in Appendix
D.
The clock error ranged from -10.9 to 10.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 05
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HRDI 01
% MLS 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/190 show the observatory data loss to be
20 hours, 45 minutes, 1.652 seconds (an increase of 9.152 seconds since
last report period). This is a 0.0348 percent data loss which equals a
99.9652 percent data capture for the mission.
The increase of 9.152 seconds occurred over two days:
* 4.928 seconds on Day 186 due to hits in the data, and
* 4.224 seconds on Day 187 due to insufficient overlap.
Testing the command side of the I/P net was conducted on 07 July and 09
July 1998:
* The 07 July testing used the UTTS and POCC AP 2. After some initial
configuration problems through NASCOM, the POCC was able to receive 32
kbps data over the I channel using the MS 422 circuit. Commanding was
also successful over the MS 422 circuit. Single commands, Ephemeris
and ATC loads were transmitted with good Spacecraft command counter
verification. 32 kbps Q channel data dumps were performed
successfully over the MS 426 line, although further testing is
required.
* The 09 July testing used the UTTS and POCC AP 5. The test was
successful. FOT transmitted ATC loads, ephemeris loads, serial
magnitude commands, and discrete commands. The ATC and Ephemeris were
dumped over the Q channel (MS 426) and compared with zero errors.
The CMS system was re-booted on 06 July 1998 due to a printer queue
anomaly. As it turns out, the CMS had not been rebooted in over 200 days,
so this latest reboot served as a good RrefreshS of the system.
The FOT generated two special MLS microprocessor loads to accommodate
testing of the MLS instrument on 10 July.
The selected date for the Orbit Adjust maneuver is under review. The
Project Manager requested a determination from the Project Scientist and
Instruments if good science could be obtained down to 580 km. This would
extend the need for an orbit adjust for a year or more. All instruments
have responded that a 580 km lower limit is acceptable for operations.
New AP ground system software testing continued. New findings in the
testing process suggest that the /DUMPCMD directive defect previously
stated already exists in the operational system.
OTHER SIGNIFICANT EVENTS
Planning continues for the UARS orbit adjust, scheduled for 12 August 1998,
and the engineering thruster tests scheduled for 28 July 1998.
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
FOT attended a meeting on 09 July 1998 to discuss the plan for testing the
UARS CMS for Year 2000 compatibility. Representatives from the FOT, MSOCC,
CMF, FDF, CMS system maintenance (hardware), CMS system administration, and
the Year 2K test team (CSC) were in attendance. The meeting was chaired by
Carey Myers and Jim Clark. Several action items were generated at the
meeting, three of which have been assigned to the FOT:
* The FOT will review the Year 2K test plan, which was provided by Sidney
Kirschner (CSC).
* The FOT will provide a list of planning aids and other files that it
receives from FDF that are necessary to facilitate thorough Year 2K
testing of the CMS.
* The FOT will attempt to provide Mr. Kirschner with any necessary CMS
related documentation.
A follow-up meeting is scheduled for 16 July 1998 to discuss the status of
the assigned action items.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 15 July 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 12 August 1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 July 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37260 187/2125 98-235 OBC Clk Adj fr 0.00 to 20.6 msec/day
37279 189/0349 98-236 OBC Clk Adj fr 20.6 to 0.00 msec/day
37288 189/1814 98-237 MLS Scan Test
37306 190/2255 98-238 HRDI To Low Power Mode
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 July 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 37217-321
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 37219-321
HRPTSPBIMON Y-HI PTS Pulse Bus Current Mon 37236-38,42
HRPSEQBIMON Y-HI PSE Quiet Bus Current Mon 37236-38,40,42,82,313
HRIFESURVTMP Y-HI IFE Survival Heater Temp. 37308
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 July 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37260 187/21:25 (06 Jul 98) from 00.0 to 20.6 msec/day
37279 189/03:49 (08 Jul 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 July 1998
03 Jul 1998 - GMT Day 184 Beta = -32.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.2 0.000 0.0
2 +11.2/ -50.4 +6.44/+4.83 +29.6/-15.2 24.8 1.021 22.5
3 +39.2/ -11.2 +5.47/+3.87 +28.8/-16.0 24.8 1.024 22.0
10 Jun 1998 - GMT Day 191 Beta = -17.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.46/-3.70 +0.0/ +0.0 18.2 0.000 0.0
2 +11.2/ -44.8 +7.09/+5.47 +29.5/-14.5 24.5 1.024 23.5
3 +33.6/ -33.6 +5.79/+4.51 +29.7/-14.9 24.5 1.023 23.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 10 July 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-044 FOT See list July 1998 14170 Generic Late Acquisition
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 10 July 1998
INSTRUMENT SUMMARY FOR MONTH OF JUNE 1998 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- ------------
ACRIM 646 74 0 0
HALOE 523 197 0 0
HRDI 239 481 0 0
MLS 062 627 0 031 Note 1
PEM (axis) 153 447 120 0
PEM (zeps) 0 720 0 0
PEM (neps) 0 720 0 0
SOLSTICE 720 0 0 0 Note 2
SUSIM 646 074 0 0
WINDII 130 590 0 0
Note 1 - MLS OFF for 31 hours due to scan mechanism anomaly
Note 2 - No star data June 27-30
Cold side of spacecraft was facing northward: 14.5 days
Cold side of spacecraft was facing southward: 15.5 days